# Learning Task-Agnostic Skill Bases to Uncover Motor Primitives in Animal Behavior

### v1

@@ 3.2 Continuous Version
Skill discovery. We model the transition kernel as an energy-based model (EBM) [8]:
[…]
where $\psi(s,a):{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}^{g}$ and $\nu(s^{\prime}):{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}^{g}$. The goal is to learn $\psi(s,a)$ and $\nu(s^{\prime})$. $Z(s,a)$ is an intractable partition function. This parametrization provides smooth, normalized probabilities with stable gradients, enabling more effective and generalizable representation learning than the unnormalized inner-product model (Eq. (3)).
Proposition 1 (Connection to Skill Definition).
Given the energy-based transition model in Eq. (7), the transition kernel can be approximated by $P(s^{\prime}|s,a)\approx\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, where $\phi(s,a)\in\mathbb{R}^{d}$ is an explicit function of $\psi(s,a)$ and the partition function $Z(s,a)$, and $\mu(s^{\prime})\in\mathbb{R}^{d}$ is a function of $\nu(s^{\prime})$.
Appendix. A contains the full proof and derivation of $\phi$ and $\mu$ in terms of $\psi$, $\nu$, and $Z$.
To learn $\psi$ and $\nu$, we sidestep the intractable computation of the partition function $Z(s,a)$ in Eq. (7) by employing NCE (Eq. (2)) to learn the representations.
@@
$s^{\prime}$ denotes a positive sample drawn from the transition distribution $P(\cdot|s,a)$, and $s^{\prime\prime}_{i}$ for $i=1,\ldots,k$ are negative samples.
Connection to behavior dynamics. With simple algebra, we obtain the quadratic potential function $P(s^{\prime}|s,a)\propto q(s^{\prime})\exp\left(\|\psi(s,a)\|^{2}/2\right)\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right)$. By enforcing unit-norm constraints $\|\psi(s,a)\|^{2}=\|\nu(s^{\prime})\|^{2}=1$, assuming $Z(s,a)$ as a constant and $q(s^{\prime})$ as uniform distribution, as well as taking $\nu$ to be the identity map, we obtain a generalized Gaussian form: $P(s^{\prime}|s,a)=\frac{1}{Z}\exp\left(-\|s^{\prime}-\psi(s,a)\|^{2}\right)$, which aligns with commonly adopted assumptions in animal behavior modeling discussed earlier. Thus, Eq. (7) offers a more general framework that extends traditional dynamics models for studying behavior. Moreover, in practice, using Eq. (7) results in a unimodal distribution over $s^{\prime}$, which closely matches the empirical structure observed in $P(s^{\prime}|s,a)$ from animal behavior data. Thus, while Eq. (3) is theoretically valid in continuous domains, we adopt the parameterization in Eq. (7) for continuous state and action spaces, as it more effectively supports the learning of skill representations.
Skill-based policy learning. After we obtain the representation $\psi(s,a)$ and $\nu(s^{\prime})$ from Eq. (8), we could theoretically get skill sets $\phi(s,a)$ expressed as the function of $\psi(s,a)$ and a normalizing term $Z(s,a)$ (see Appendix. A). However, since in practice $Z(s,a)$ remains intractable, even with the optimal $\psi$ it is still hard to obtain $\phi$ exactly. Therefore, we introduce a mapping $f:\psi\rightarrow\phi$, parameterized with a lightweight neural network, and learn it via policy learning. Our aim is to learn a function $f$ so that $\phi=f(\psi)$ yields optimal basis functions of policy that best account for the animal behavior data. By applying $\phi=f(\psi)$ to Eq. (3), we obtain $Q(s,a)=f\bigl{(}\psi(s,a)\bigr{)}^{\top}u$.
As mentioned earlier, learning both $f$ and $u$ using the MLE objective in Eq. (6) becomes intractable for continuous data, as the denominator involves integration over an unbounded continuous action space. This brings us back to the challenge of estimating an unnormalized energy function, $\pi(a|s)\propto\exp(Q(s,a))\propto\exp(f(\psi(s,a))^{\top}u)$. Thus, it’s reasonable to apply NCE here again,
@@
3.3 Understanding Animal Behavior via Skill and Policy Learning
In this section, we discuss how motor skills $\phi$ and skill weights $u$ can be used to describe animal behavior trajectories, particularly allowing $u(t)$ to vary across tasks or time points. These time-varying coefficients dynamically modulate the influence of each skill. We consider two behavioral modeling scenarios: (1) discrete state-action spaces in a multi-task setting, and (2) continuous state-action spaces in a time-varying setting.
Scenario (1): Consider a mouse navigating a maze [23], where trajectories are discretized into a finite state space (locations) and actions are discrete (up, down, left, right, stay). We assume the animal switches between different reward functions to guide its navigation, with the timing of each reward condition known [24]. We first learn shared skills $\phi(s,a)$ using Eq. (3), then fit task-specific policies $\pi(a|s,t)$ using Eq. (5), where each task $t$ corresponds to one of $T$ reward conditions. This yields $T$ sets of weights $u(t)$, one per task, while sharing a common skill representation across tasks.
Scenario (2): A representative case is a freely behaving mouse [1, 2], where the state is defined by pose keypoints and the action by state change—both continuous. We first learn $\psi(s,a)$ from pose trajectories using Eq. (8). Assuming the policy evolves smoothly over time, we learn $u(t)$ and $f$ via Eq. (9), yielding the motor skill $\phi(s,a)=f(\psi(s,a))$ and time-varying policy $\pi(a|s,t)$. Unlike models with abrupt discrete switches, this continuous-time formulation captures gradual behavioral changes more faithfully over long pose sequences.
Beyond capturing smoothly time-varying skill compositions, our framework allows multiple skills to be active simultaneously. Since the policy is defined as $\pi(a|s,t)\propto\exp\bigl{(}\phi(s,a)^{\top}u(t)\bigr{)}$, each action is a linear combination of basis skills weighted by $u(t)$, enabling overlapping and composable behaviors. For instance, back grooming may blend grooming and turning back, while side grooming mixes grooming with turning side—executed concurrently by different body parts. Unlike discrete switching-state models, which assign one behavior per state, our continuous skill-based approach provides a more flexible and interpretable representation of complex pose dynamics and, to our knowledge, is the first to offer a fully compositional and continuously time-varying description of animal trajectories.
@@ 3.4 Reward Recovery
After estimating $u(t)$ as the skill weights, we can further infer $w(t)$ in the linearized reward model $r(s,a,t)=\phi(s,a)^{\top}w(t)$ as $w(t)=u(t)-\gamma\int V(s^{\prime},t)\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$, where $V(s,t)=\log\sum_{a}\exp Q(s,a,t)$. This allows us to recover the time-varying reward function $r(s,a,t)$ used by animals. Recovering the internal reward function aligns with the goals of inverse reinforcement learning (IRL) [25], where both the policy and underlying reward are inferred. In the context of animal behavior [24, 26, 27], identifying such rewards offers insight into the internal motivations driving behavior and provides a window into animal cognition and decision-making processes. Since $V(s,t)$ can be easily computed in closed form in discrete settings, we validate our method by visualizing the inferred rewards in the first two experiments, where the state-action space is finite.
4 Related work
Skill learning from human behavior. Recent work on skill learning for human behavior understanding has primarily focused on robotics, using deep RL and imitation learning to extract reusable primitives from demonstrations [28, 29, 30, 31]. Parallel advances in IRL aim to improve policy inference using low-dimensional bases [32]. However, these approaches are designed for artificial agents in structured settings and do not capture the fine-grained, time-varying structure of naturalistic behavior. In contrast, our work uses RL-based representation learning on real animal data to uncover task-agnostic motor skills and learn a skill-based policy that models continuous, compositional behavior dynamics—capabilities absent in prior work.
Motor primitives learning from animal behavior. Most existing methods for understanding animal behavior identify motor primitives using classification [33, 34], clustering [4, 5], or behavior segmentation techniques [1, 2, 3]. Existing methods label each time point with discrete syllables (e.g., walking, running) and analyze neural correlates of syllable transitions or structure, but treat behavior as rigid and predefined. They fail to capture its compositional, continuously time-varying nature. To our knowledge, no prior work models motor primitives from this flexible perspective. Our approach introduces representation and imitation learning into animal behavior research, enabling dynamic and interpretable modeling of behavior and internal motivation.
@@ 5.1 Application to simulated data in a multi-task gridworld
We evaluate our algorithm on a multi-task gridworld environment. The gridworld consists of a $3\times 3$ lattice with nine discrete states, and each state allows four possible actions: Up, Down, Left, and Right (Fig. 2A). In task $i$, a high reward is assigned to the $(s,a)$ pairs that move toward the location $i$. Fig. 2C (left) shows the ground truth of reward functions for all nine tasks. In each episode, the agent starts from a random start state and must navigate to the task-specific location $i$.
@@
As described in Sec. 3.1, we then learn a set of latent skills and use them to construct the task-specific policy $\pi(a|s,t)$ for each task $t\in\{1,...,9\}$. We assume a total of 64 skills and learn the 64 $\phi$ vectors. Visualizations of the raw features for these skills are shown in the Appendix. B. Using the learned skills, we recover the policy and further infer the reward function used by the agent, as described in Sec. 3.4, with the form: $r(s,a,t)=\phi(s,a)^{\top}w(t)$. This recovered reward (Fig. 2C, right) closely matches the ground truth, achieving a Pearson correlation coefficient of 0.96, indicating that the learned skills are sufficient for accurately reconstructing the reward function from behavior data.
To better interpret the learned skills and their role in reward composition, we apply principal component analysis (PCA) to the $\phi$ matrix and find that only 8 principal components capture most of the variance (Fig. 2B). Therefore, we visualize the top 8 PC skills and their corresponding task-specific coefficients $w(t)$ in Fig. 2D, treating them as basis vectors spanning the skill space.
@@
The PC skills exhibit interpretable patterns. For instance, in skill 0, $(s,a)$ pairs leading to the bottom-left grid have strong positive values, while those leading to the middle-right grid have strong negative values. This skill corresponds to moving away from the middle-right grid toward the bottom-left. Examining the $w$ matrix, we see that skill 1 has a strong positive weight for task 8, consistent with the goal of moving toward the bottom-right corner in that task. In contrast, skill 0 contributes negatively to task 8, as it promotes movement toward the bottom-left and away from the goal. Similar interpretations can be made for other skills and tasks.
@@
Overall, in this gridworld example, we successfully recover reward functions from behavior trajectories, and importantly, the learned $\phi$ and $w$ are both interpretable and aligned with task goals.
@@
Figure 2: A. State-action map: each of the nine grids is divided into four cells representing action values (Up, Down, Left, Right). In task $i$, high reward is assigned to $(s,a)$ pairs moving toward the $i$th location. B. Explained variance of the top 15 principal components; variance drops near zero after PC7. C. Left: true rewards for all 9 tasks (yellow = high, green = low). Right: recovered rewards. D. Reward weight $w$ and top 8 PC skills from the $\phi$ matrix. Reward weights indicate the contribution of the top 8 PC skills to each task. In the PC skill plot, red indicates positive feature values, blue indicates negative.
@@ 5.2 Application to animal navigation behavior
5.2.1 Dataset and model setup
@@
We next evaluate our algorithm on real animal behavior data from [23]. In this experiment, a thirsty mouse is trained to navigate in a maze (Fig. 3A), starting each trial from the central home cage and attempting to reach a water port located at the end of one branch. Although the mouse’s behavior is primarily driven by water foraging after training, it also exhibits exploration of unvisited areas and returns to the home cage for shelter. This complex behavior cannot be captured by a single reward function. The state space has the structure of a binary tree and is defined by the mouse’s location. The actions include moving to its left parent, right parent, left child, and right child. Non-existence of the moving target will end in staying still.
@@
We first apply the EM-based segmentation algorithm from [24] to divide the behavior into three interpretable tasks: water seeking, home seeking, and exploration. The algorithm could also infer reward maps for these tasks, which are shown in Fig. 3A (top row). While the mouse’s true internal reward functions remain unknown, we treat these inferred rewards as effective ground truth since they generate policies that closely replicate the observed behavior. Thus, we are able to segment long navigation trajectories into task-specific episodes, yielding a multi-task dataset with discrete $(s,a)$ pairs, analogous to our simulated gridworld setting.
To uncover a task-agnostic set of skills capable of constructing all three reward functions, we apply our model as described in Sec. 3.1. The learned $\phi$ vectors serve as a policy basis, encoding movement tendencies such as approaching or avoiding certain regions of the maze.
@@
Figure 3: A. True and recovered rewards for the three tasks. B. Two dominant skills for the water and home tasks respectively. Each skill indicates that taking a specific action (orange arrow) toward the dark purple state yields a high value. C. $w$ values for all tasks, colored by task; blue stars highlight high-weight skills for water seeking, and orange for home seeking.
5.2.2 Results

### v2

@@ 3.2 Continuous Version
While learning from discrete data is relatively straightforward, the continuous case presents additional challenges for two main reasons. First, in the motif discovery step (Eq. 2), the decomposition $P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$ is too restrictive to capture the complexity of continuous behavioral dynamics, such as pose transitions in freely moving animals (Weinreb et al., 2024). For example, consider a common and biologically plausible behavioral model: $s^{\prime}=h(s,a)+\epsilon$, where $h$ is a dynamics function and $\epsilon$ is Gaussian noise. This additive structure, widely used in behavioral modeling, contrasts with the multiplicative form $\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, suggesting that parameterizations preserving additive relationships between $\{s,a\}$ and $s^{\prime}$ are more suitable. Second, in motif-based policy learning, for discrete datasets with a small action space, the denominator (partition function) in Eq. 5 is easy to compute. But for continuous data, the action space is infinite, making it infeasible to enumerate all actions and integrate. In light of these challenges, we adopt an alternative approach to learn the motif representations and policy in continuous state-action spaces.
Motif discovery. We model $P(s^{\prime}|s,a)$ as an energy-based model (EBM) (Shribak et al., 2024):
$P(s^{\prime}|s,a)=q(s^{\prime})\exp\left(\psi(s,a)^{\top}\nu(s^{\prime})-\log Z(s,a)\right),Z(s,a)=\int q(s^{\prime})\exp(\psi(s,a)^{\top}\nu(s^{\prime}))ds^{\prime},$ (6)
where $\psi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{g}$ and $\nu:\mathcal{S}\to\mathbb{R}^{g}$ are neural-network feature maps. Here, $Z(s,a)$ is an intractable partition function. Compared to the unnormalized inner-product model (Eq. 2), this EBM formulation yields smooth, normalized probabilities and stable gradients, leading to more effective and generalizable motif representations.
Proposition 2 (Connection to Motif Definition).
Given the EBM model in Eq. 6, the transition kernel can be approximated by $P(s^{\prime}|s,a)\approx\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, where $\phi(s,a)\in\mathbb{R}^{d}$ is an explicit function of $\psi(s,a)$ and the partition function $Z(s,a)$, and $\mu(s^{\prime})\in\mathbb{R}^{d}$ is a function of $\nu(s^{\prime})$.
Appendix. C contains the full proof and derivation of $\phi$ and $\mu$ in terms of $\psi$, $\nu$, and $Z$.
To learn $\psi$ and $\nu$, we employ noise-contrastive estimation (NCE) (Ma & Collins, 2018; Gutmann & Hyvärinen, 2010; 2012), which enables optimization of unnormalized statistical models without explicitly computing the partition function. In this way, we sidestep the intractable computation of $Z(s,a)$ in Eq. 6 by solving
@@
where $s^{\prime}$ denotes a positive sample drawn from the transition distribution $P(\cdot|s,a)$, and $s^{\prime\prime}_{i}$ for $i=1,\ldots,k$ are negative samples from distribution $\rho$.
Connection to behavioral dynamics. With simple algebra, we obtain the quadratic potential function $P(s^{\prime}|s,a)\propto q(s^{\prime})\exp\left(\|\psi(s,a)\|^{2}/2\right)\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right)$ from Eq. 6. By enforcing unit-norm constraints $\|\psi(s,a)\|^{2}=\|\nu(s^{\prime})\|^{2}=1$, assuming $Z(s,a)$ as a constant and $q(s^{\prime})$ as uniform distribution, as well as taking $\nu$ to be the identity map, we obtain a generalized Gaussian form: $P(s^{\prime}|s,a)=\frac{1}{Z}\exp\left(-\|s^{\prime}-\psi(s,a)\|^{2}\right)$, which aligns with commonly adopted assumptions in animal behavior modeling discussed earlier. Thus, Eq. 6 offers a more general framework that extends traditional dynamics models for studying behavior. Moreover, in practice, using Eq. 6 results in a unimodal distribution over $s^{\prime}$, which closely matches the empirical structure observed in $P(s^{\prime}|s,a)$ from animal behavior data. Thus, while Eq. 2 is theoretically valid in continuous domains, we adopt the parameterization in Eq. 6 for continuous state and action spaces, as it more effectively supports the learning of motif representations underlying animal behavior.
Motif-based policy learning. After we obtain the representation $\psi(s,a)$ and $\nu(s^{\prime})$ from Eq. 7, we could theoretically get motif sets $\phi(s,a)$ expressed as the function of $\psi(s,a)$ and a normalizing term $Z(s,a)$ (see Appendix. C). However, since in practice $Z(s,a)$ remains intractable, even with the optimal $\psi$ it is still hard to obtain $\phi$ exactly. Therefore, we introduce a mapping $f:\psi\rightarrow\phi$, parameterized with a neural network, and learn it via policy learning. Our aim is to learn a function $f$ so that $\phi=f(\psi)$ yields optimal basis functions of policy that best account for the animal behavior data. By applying $\phi=f(\psi)$ to Eq. 3, we obtain $Q(s,a)=f\bigl(\psi(s,a)\bigr)^{\top}u$.
As mentioned earlier, learning both $f$ and $u$ using the MLE objective in Eq. 5 becomes intractable for continuous data, as the denominator involves integration over an unbounded continuous action space. This brings us back to the challenge of estimating an unnormalized energy function, $\pi(a|s)\propto\exp(Q(s,a))\propto\exp(f(\psi(s,a))^{\top}u)$. Thus, it’s reasonable to apply NCE here again,
@@
3.3 Understanding Animal Behavior via Motif and Policy Learning
In this section, we discuss how motor motifs $\phi$ and motif weights $u$ can be used to describe animal behavior trajectories, particularly allowing $u$ to vary across tasks $t$ or time points $t$. As the motif coefficients, $u(t)$ would dynamically modulate the influence of each motif on the final policy. We consider two behavioral modeling scenarios: (1) discrete state-action spaces in a multi-task setting, and (2) continuous state-action spaces in a time-varying setting. The concrete results will be later shown in Sec. 4.2 and Sec. 4.3 respectively. In either case, $u(t)$ would be a matrix where each column corresponds to one weight for one task or time point.
Scenario (1): Consider a mouse navigating a maze (Rosenberg et al., 2021), where trajectories are discretized into a finite state space (locations) and actions are discrete (up, down, left, right, stay). We assume the animal switches between $T$ strategies, each associated with a unique reward, that guide its navigation, with the timing of each reward condition known from Ke et al. (2025). We first learn shared motifs $\phi(s,a)$ using Eq. 2, then fit task-specific policies $\pi(a|s,t)$ using Eq. 5, where each task $t$ corresponds to one of $T$ strategies. This yields $T$ sets of weights $u(t)$, one per task, while sharing a common motif set across tasks.
Scenario (2): A representative case is a freely behaving mouse (Wiltschko et al., 2015; Weinreb et al., 2024), where the state is defined by pose keypoints and the action by state change–both continuous. We first learn $\psi(s,a)$ from pose trajectories using Eq. 7. Assuming the policy evolves smoothly over time, we learn $u(t)$ and $f$ via Eq. 8, yielding the motor motif $\phi(s,a)=f(\psi(s,a))$ and time-varying policy $\pi(a|s,t)$. Unlike models with abrupt discrete switches, this continuous-time formulation captures gradual behavioral changes more faithfully over long pose sequences.
Beyond capturing smoothly time-varying motif compositions, our framework allows multiple motifs to be active simultaneously. Since the policy is defined as $\pi(a|s,t)\propto\exp\bigl(\phi(s,a)^{\top}u(t)\bigr)$, each action is a generalized linear combination of basis motifs weighted by $u(t)$, enabling overlapping and composable behaviors. For instance, back grooming may blend grooming and turning back, while side grooming mixes grooming with turning side—recruiting different motifs concurrently. Unlike discrete switching-state models, which assign one behavior per state, our continuous motif-based approach provides a more flexible and interpretable representation of complex pose dynamics and, to our knowledge, is the first to offer a fully compositional and continuously time-varying description of animal trajectories. In Sec. 4, we would show (1) what motifs we have learned, and (2) how they are used to construct the final policy, on one simulation datasets and two real animal behavior datasets.
@@ 3.4 Reward Recovery
After estimating $u(t)$ as the motif weights for policy construction, we can further infer $w(t)$ for reward representation $r(s,a,t)=\phi(s,a)^{\top}w(t)$ as $w(t)=u(t)-\gamma\int V(s^{\prime},t)\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$, where $V(s,t)=\log\sum_{a}\exp Q(s,a,t)$. This allows us to recover the time-varying reward function $r(s,a,t)$ used by animals. Recovering the internal reward function aligns with the goals of inverse reinforcement learning (IRL) (Ziebart et al., 2008), where both the policy and underlying reward are inferred from demonstrations. In the context of animal behavior (Ke et al., 2025; Zhu et al., 2024; Ashwood et al., 2022), identifying such rewards offers insight into the internal motivations driving behavior and provides a window into animal cognition and decision-making processes. Since $V(s,t)$ can only be easily computed in closed form in discrete settings, we validate our method by visualizing the inferred rewards in the first two experiments, where the state-action space is discrete and finite.
@@ 4.1 Application to simulated data in a multi-task gridworld
The gridworld consists of a $3\times 3$ lattice with nine discrete states, and each state allows four possible actions: Up, Down, Left, and Right (Fig. 2A). In task $i$, a high reward is assigned to the $(s,a)$ pairs that move toward the location $i$. Fig. 2C (left) shows the ground truth of reward functions for all nine tasks. In each episode, the agent starts from a random start state and must navigate to the task-specific location $i$. See Appendix. D for more details for data generation.
@@
As described in Sec. 3.1, we learn a set of latent motifs and use them to construct the task-specific policy $\pi(a|s,t)$ for each task $t\in\{1,...,9\}$. Because the computational complexity only scales linearly with the number of motifs, we assume a total of 64 motifs and learn the 64 $\phi$ vectors to cover the motif space as much as possible. Visualizations of these motifs are shown in the Appendix. D. Using the learned motifs, we recover the policy and further infer the reward function, as described in Sec. 3.4, with the form: $r(s,a,t)=\phi(s,a)^{\top}w(t)$. This recovered reward (Fig. 2C, right) closely matches the ground truth, achieving a Pearson correlation coefficient of 0.96, indicating that the learned motifs are sufficient for accurately reconstructing the reward function from behavior data.
@@
To better interpret the learned motifs and their role in reward composition, we apply principal component analysis (PCA) to the $\phi$ matrix and find that only 8 principal components capture most of the variance (Fig. 2B). Therefore, we visualize the top 8 PC motifs and their corresponding task-specific coefficients $w(t)$ in Fig. 2D, treating them as basis vectors spanning the motif space. The PC motifs exhibit interpretable patterns. For instance, in motif 0, $(s,a)$ pairs leading to the bottom-left grid have strong positive values, while those leading to the middle-right grid have strong negative values. This motif corresponds to moving away from the middle-right grid toward the bottom-left. Examining the $w$ matrix, we see that motif 1 has a strong positive weight for task 8, consistent with the goal of moving toward the bottom-right corner in that task. In contrast, motif 0 contributes negatively to task 8, as it promotes movement toward the bottom-left and away from the goal. Similar interpretations can be made for other motifs and tasks.
@@
In this gridworld experiment, we successfully recover reward functions from behavior trajectories, and, importantly, the learned $\phi$ and $w$ are effectively deployed in different task settings.
@@
Figure 2: A. State-action map: each of the nine grids is divided into four cells representing action values (Up, Down, Left, Right). In task $i$, high reward is assigned to $(s,a)$ pairs moving toward the $i$th location. B. Explained variance of the top 15 principal components; variance drops near zero after PC7. C. Left: true rewards for all 9 tasks (yellow = high, green = low). Right: recovered rewards. D. Reward weight $w$ and top 8 PC motifs from the $\phi$ matrix. Reward weights indicate the contribution of the top 8 PC motifs to each task. In the PC motif plot, red indicates positive feature values, blue indicates negative.
@@ 4.2 Application to animal navigation behavior
Dataset and model setup.
@@
We next evaluate our method on a real animal behavior dataset from Rosenberg et al. (2021). In this experiment, a thirsty mouse is trained to navigate in a binary-tree maze (Fig. 3A), starting each trial from the central home cage and attempting to reach a water port at one leaf node. The state space is defined by the mouse’s location on the tree. The actions include moving to its left parent, right parent, left child, and right child. Although the mouse’s behavior is primarily driven by water foraging, it also exhibits exploration of unvisited areas and returns to the home cage for shelter. This complex behavior cannot be captured by a single reward function. Studying this dataset allows us to discover motifs shared across multiple reward functions and policies. This in turn tests whether complex navigation behavior, under multiple competing motivations, can be distilled into a small set of reusable motifs that provide interpretable insight into animal decision-making.
@@
We first apply the segmentation algorithm from Ke et al. (2025) to divide long behavior trajectories into three interpretable tasks: water seeking, home seeking, and exploration. The algorithm could also infer reward maps for these tasks, which are shown in Fig. 3A (top row). While the mouse’s true internal reward functions remain unknown, we treat these inferred rewards as effective ground truth since they generate policies that closely replicate the observed behavior. Thus, we are able to segment long navigation trajectories into task-specific episodes, yielding a multi-task dataset with discrete $(s,a)$ pairs, analogous to our simulated gridworld setting. To uncover a task-agnostic set of motifs capable of constructing all three reward functions, we apply our model as described in Sec. 3.1. Again, to cover the motif space as much as possible, we set the number of motifs $d=128$.
@@
Figure 3: A. True and recovered rewards for the three tasks. B. Two dominant motifs for the water and home tasks respectively. Each motif indicates that taking a specific action (orange arrow) toward the dark purple state yields a high value. C. $w$ values for all tasks, colored by task; blue stars highlight high-weight motifs for water seeking, and orange for home seeking, all shown in (B).
Results.